### v2

# Ina K. Schieferdecker Technische Universität Berlin, Einsteinufer 25, 10587 Berlin, Germany

Abstract
The effectiveness of model-driven software engineering (MDSE) has been successfully demonstrated in the context of complex software; however, it has not been widely adopted due to the requisite efforts associated with model development and maintenance, as well as the specific modelling competencies required for MDSE. Concurrently, artificial intelligence (AI) methods, particularly deep learning methods, have demonstrated considerable abilities when applied to the huge code bases accessible on open-source coding platforms. The so-called big code provides the basis for significant advances in empirical software engineering, as well as in the automation of coding processes and improvements in software quality with the use of AI. The objective of this paper is to facilitate a synthesis between these two significant domains of software engineering (SE), namely models and AI in SE. The paper provides an overview of the current state of AI-augmented software engineering and develops a corresponding taxonomy, AI4SE. In light of the aforementioned considerations, a vision of AI-assisted Big Models in SE is put forth, with the aim of capitalising on the advantages inherent to both approaches in the context of software development. Finally, the new paradigm of pair modelling in MDSE is proposed.
keywords:
Software Engineering, Model-Driven Engineering, Modelling, Domain Modelling, Artificial Intelligence, Big Models, AI4SE
1 Introduction
"Programming at the beginning of 2025 and at the end of 2025 are going to be completely different things."
Sam Altman, Febr. 7, 2025, The Age of AI – Panel Discussion, TU Berlin, Germany
The advent of artificial intelligence (AI), and more specifically, generative AI, has precipitated a substantial transformation in the field of software engineering (SE), which is expected to undergo further significant changes in the near future (see also the quotation by Sam Altman, OpenAI). In order to better understand and comprehend these developments, this paper analysis the state of art in using AI for SE Software engineering is a field of informatics/computer science that addresses the development and analysis of systematic approaches for the design, development, verification & validation and maintenance of software and of software-based systems [^1] In this paper, the term ”software” will be used as a general reference to the collective term for computer programs and related SE artefacts.. It establishes principles and identifies optimal practices for the production and operation of software, with the aim of ensuring the reliability, security, scalability, maintainability and alignment of software with user, business and societal requirements. The objective of software engineering is the production of high-quality software that is cost-effective, delivered in a timely manner and can be readily adapted as requirements evolve.
The term software engineering was coined in 1968 with the very first software crisis [82], but still software engineering practices and the resulting software quality have not kept pace with the quality levels required in critical domains or application contexts. In 2024, another dramatic story was added to the software horror show: The millions of enterprise outages caused worldwide by the CrowdStrike Falcon sensor update [24] were due to simple errors such as not validating the number of fields in an IPC template type, missing a runtime array bounds check, a logic error in the content validator, not testing template types enough, and not testing template instances within the content interpreter. This was unprofessional software development that ignored state of the art of SE.
However, software outages are not the only indication that SE still has no full answers to the challenges posed by the increasing complexity of software-based systems and the diversity of requirements placed on them: [71] analysed that "[t]here exists a statistically significant medium sized difference between open and closed source projects: the former have a DD [(defect density)] that is 4 defects per KLoC lower than the latter. Java projects exhibit a significantly lower DD than C projects, 4.1 defects per KLoC on average: In general the [code] size appears to be negatively correlated to DD: the larger the project the lower the DD. In particular, large projects are 10 times less defective than medium ones.". According to [59], the defect density (DD) of software projects within the industrial sector is estimated to fall within the range of 1 to 25. This suggests that a software program comprising one million lines of code (MLoC) may contain between 1,000 and 25,000 defects. Given the lines of code in a piece of software, deriving estimates of the approximate number of defects in it is a relatively straightforward process. To illustrate, [74] compares typical software such as operating systems, browsers, and office suites: The Windows 10 operating system has approximately 80 MLoC, Ubuntu 50 MLoC, MacOS X 84 MLoC, Android 12 MLoC, iOS 12 MLoC, the browser Google Chrome 6.7 MLoC, Mozilla Firefox 21 MLoC, or the office suites Microsoft Office 2013 45 MLoC, Apache OpenOffice 19 MLoC, and LibreOffice 10 MLoC. Therefore, the challenges associated with the development of accurate and suitable software remain significant.
[17] argued that the primary challenge in software development lies in the specification, design, and testing of complex systems, rather than in the labour for representation of the system or testing its fidelity. It emphasized the distinction between essential difficulties (inherent complexity) and accidental difficulties (extraneous challenges), stating that past advances, like high-level programming languages, have only reduced the latter. The paper suggested that addressing accidental difficulties would not lead to major improvements in software development, as essential difficulties remain fundamental. [36] critiqued the notion of "accidental difficulties", arguing that these so-called accidents are often the result of negligence or poor practices, not mere happenstance. This paper advocated for a disciplined, science-based, and model-driven approach to software development, similar to traditional engineering disciplines. Building on this, [68] highlighted the growing societal dependence on autonomous and intelligent software systems, suggesting that new approaches are needed to ensure not only traditional software quality facets (safety, security, etc.), but also to address socio-technical and socio-political implications, particularly in human-machine collaboration.
In model-driven development (MDD [70]), also known as model-driven software engineering (MDSE [69]), models are original artefacts that are engineered with the intention of facilitating the top-down construction of complex software. Furthermore, models are utilised during runtime to enable the monitoring, verification & validation of software operations, as well as the optimisation of its performance [41, 34, 85]. The issues in using traditional (top-down) MDSE has been e.g. analysed in [35].
As has been the case with the comprehensive incorporation of modelling in software engineering, AI techniques have been employed in software engineering activities from the outset, see for example [11]. In the context of significant advancements in AI-driven software engineering in 2021, Gartner’s hype cycle for emerging technologies [19] projected that AI-enhanced software engineering would reach the peak of inflated expectations within a five-to-ten-year timeframe, subsequently entering a phase of productivity stabilization. Since 2023, it is estimated that AI-augmented software engineering will reach the productivity plateau already in two to five years, while in previous iterations of the hype cycle, AI-augmented development and AI-assisted design were identified as emerging technologies on the rise. It was interesting to observe the evolving business expectations and the convergence of trends for AI-augmented design and development into a unified technological trend for AI-augmented engineering. Moreover, it is encouraging to observe that software engineering is receiving particular attention and prioritisation, which is indicative of the growing importance of this field.
Therefore, this paper focuses on the intersection of AI and SE, specifically MDSE, to improve software engineering practices. The paper aims to bridge the gap between the theoretical advantages of MDSE and its practical limitations by integrating AI techniques, and introduces a novel taxonomy, AI4SE, to classify AI applications within the software engineering domain, highlight open research questions, and direct future research. The paper uses a combination of literature review, analysis of existing software engineering practices and AI applications, and development of a novel taxonomy to achieve its objectives. The main contributions of the paper include:
  • The new AI in Software Engineering Taxonomy AI4SE. This taxonomy categorises AI applications based on the purpose of using AI, the target of AI within software engineering (e.g. requirements, design, coding, testing, operations, process), the type of AI used, and the level of autonomy achieved by the AI.
  • The novel Modelling Hypothesis, which posits that models have statistical properties that can be used to develop more effective software engineering tools.
  • The concept of Big Models, analogous to big code, as a basis for advanced empirical MDSE. The paper suggests that these models can be generated top-down, or extracted bottom-up from code, and that AI methods can be applied to these big models.
  • The new paradigm of pair modelling in MDSE, where a software engineer and an AI tool/agent collaborate in software development, similar to pair programming. The aim is to leverage the strengths of both human and AI for MDSE to improve software quality.
The AI4SE taxonomy is validated by categorising highly cited and recent publications in AI-assisted software engineering. This classification shows the various ways AI is being applied across different activities in SE, such as requirements engineering, architecture design, coding, testing, and deployment.
In section 2, the paper addresses the role of models in software engineering. The evolution of big code on coding platforms is discussed in section 3 and the utilisation of AI for coding and software engineering in section 4. The novel taxonomy of AI for SE (AI4SE) to facilitate a deeper understanding of this evolving field is presented also in this section. Section 5 examines the practice of contributing software models to coding platforms, which is increasingly being done together with software code. In line with the increasing prevalence of large-scale models, it examines recent developments in AI-enhanced modelling and model utilisation within the domain of software engineering and presents the Big Model concept and the pair modelling paradigm. In section 6, related work is reviewed. A summary concludes the paper.
2 Models in Software Engineering
According to the IEEE Software Engineering Body of Knowledge [73] (see Figure 1), the software lifecycle is comprised of distinct phases and activities that can be described as knowledge areas, including requirements, architecture, design, construction, testing and maintenance. There are also knowledge areas that deal with the fundamentals of computer science, mathematics and engineering, as well as cross-cutting activities pertaining to software quality, security, configuration management and engineering management. Furthermore, there are cross-cutting knowledge areas that encompass the processes, operations, professional practice and economics of software engineering, along with a distinct area for software engineering models and methods. This knowledge area encompasses the processes and methodologies associated with modelling, the various types of models, including those pertaining to information, behaviour and structure, and the analysis of models, see also [69]. Overall, SWEBOK addresses dedicated model-based or model-driven approaches in each knowledge area, such as model-based requirements specification, model-based architecture, model-driven design, lifecycle models, executable models, model-based testing, etc.
Figure 1: Software Engineering knowledge areas according to SWEBOK [73]
This is because models in SE are the key to dealing with the complexity of software: Models provide the essential abstractions to capture requirements, support design decisions, or to offer comprehensive overviews of software structures and behaviours. They are essential tools in the engineering process for constructing and maintaining software. They are also vital for configuration, monitoring and other runtime support during software operation and management. The understanding of Model-Driven Software Engineering (MDSE) has evolved in recent years as the capabilities of computing, modelling and the available software and data have evolved:
  1. 1.
    In addition to explicit model development and top-down model use, models can also be extracted from software executions, from software monitorings or from code artefacts with the objective of avoiding the costs associated with model development on the one hand and leveraging the benefits of MDSE on the other hand [7].
  2. 2.
    The combination of model development of model artefacts with model extraction from code artefacts has resulted in an hybrid top-down/bottom-up MDSE approaches [76] benefiting from both directions of MDSE as shown in [83, 85] or [37].
  3. 3.
    Furthermore, as the abstractions inherent in programming became more intricate, low-level information and structure models have become indispensable to conventional programming. This phenomenon is exemplified by the utilisation of data structures that are frequently defined in JSON, XML schemata, or SQL table specifications, to name but a few.
These practical outcomes of model-driven, model-based and/or model-oriented software developments established a foundation for the utilisation of AI in software engineering, as discussed in Section 5. Before that, however, section 3 will take a closer look at the current state of big code on open source platforms, and section 4 will take a closer look at the state of applying AI to big code.
3 Big Code in Software Engineering
The advent of software coding platforms can be traced back to the 1990s, although it was not until the early 2000s that they truly gained prominence, with the launch of GitHub in April 2008 representing a significant milestone in this regard. At the present time, GitHub is the most widely utilised software coding platform. As stated by [15], GitHub hosts over 100 million projects and 40 million users. It became not only a significant platform for software engineering collaboration, but also a prominent reference for open-source software mining [45]. The study also demonstrated that a considerable number of GitHub repositories are not directly related to software development. This is because GitHub is not solely utilised for coding collaborations; it is also employed for collaborations on websites, editing books or other publications, and is even used as a storage platform. A later manual analysis by the same authors [46] revealed that over a third of the repositories on GitHub were not software development repositories, which degraded now to roughly 20 percent according to [4] (see figure 2 upper left).
Nevertheless, the software engineering community can further improve its understanding of software engineering processes, methods and tools by gaining insights through the mining of coding platforms using tools such as Kibble [4]. In a Feb. 2025 Kibble snapshot analysing lines of code from all GitHub code repositories, about 70% is actual code, while the other 30% is comments and empty lines. Java is the most common coding language, followed by XML and JSON for data and data structures (see figure 2 below left). A GitHub search counting files in GitHub repositories shows the ratio of the number of files for coding files to model files for software designs, interface models and data models (see figure 2 top right). The split between design models, interface models and data models is shown in figure 2 below right, which shows that about 1% of the files on GitHub contain models or partial model information.
Figure 2: Repositories breakdown by Kibble and GitHub Search taken 4-6 February 2025
All of these figures are only rough indicators: Lines of code and number of files are crude measures of software. File extensions do not accurately indicate file types. Coding and modelling tools are numerous and sometimes have proprietary file extensions. Overall, a GitHub search for more specific file types such as UML works better than for more general file types such as XML or JSON, but can still lead to false positives.
Nevertheless, the sheer numbers show that modelling is being used in software development. A more detailed approach to modelling on GitHub, including image analysis, is given in [72], but this is not applicable to GitHub as a whole due to API limitations. Despite this, the numbers show that modelling is being actively used. In fact, modelling is mainly used for data/interface structures, due to the simple fact that data and interfaces do not work without at least a well-defined structure. In contrast, design and behaviour models for complete software (packages) are only used for a small fraction.
This is very much in line with the analysis of [34] over a decade ago: Code generators may produce a large amount of code, but not all of it is necessarily useful or would have been written manually. The effort required to develop, maintain and adapt the generators reduces the productivity gains from modelling. When manual code changes are required, they make round-trip engineering difficult. Generated code must be checked by developers to ensure correctness, limiting the benefits of automation. Poor software design models can lead to inefficient development. As a result, developers often prefer generic solutions with hand-written, reusable code to customised, generated solutions. As programming languages and frameworks become more expressive, the need for model-driven code generation diminishes. In this sense, GenAI further empowers the software developer. But does this mean that MDSE will disappear, or will the ability to (re)generate models from code and/or keep them in sync with software requirements evolution and manual code changes result in a boost for MDSE (see section 5)?
Back to big code: Due to the large number of software repositories with roughly 500 MB code each, the concept of "Big Code", analogous to the notion of "Big Data", was introduced in reference to the vast quantities of software code that have accumulated over time [58, 61, 84]. This code can be reused, extended and analysed to gain deeper insights into software engineering. It can be used to train ML models and enabled the application of AI for SE in large scale [53].
4 Artificial Intelligence in Software Engineering
The use cases for AI in SE are numerous and cover a wide spectrum as discussed already back in 1988 [11] and since then [2]. For example, [31] proposes the AI in SE Application Levels (AI-SEAL) taxonomy, which differentiates between the point of applying AI to the software engineering process, the software product or at runtime, the levels of automation, and the types of AI along the five tribes differentiation by [27]. It is noteworthy that the majority of the papers analysed in [31] employ AI in the software engineering process. However, despite the numerous facets of the software engineering process (Figure 1), this study does not provide further elaboration. Nevertheless, other publications offer more detailed discussions of AI in SE, including [10] or [62]. Notwithstanding the aforementioned considerations, a taxonomy for the role of AI in software engineering has yet to emerge.
In light of the recent advancements in machine learning (ML) that have made significant breakthroughs possible, this paper focuses mainly on the latest developments in the application of ML to SE. For ML to be effective, it is essential to utilise the appropriate structures inherent to SE artefacts, which has been a topic of considerable debate: [3] presents an overview of the various ways in which source code can be represented, including representational models of tokens, token contexts, program dependency graphs, API calls, abstract syntax trees, object usage, and others. These representational models are used for AI assistance in SE, including the creation of recommender systems, the inference of coding conventions, the detection of anomalies and defects, the analysis of code, the rewriting and translation of code, the conversion of code to text for the purposes of documentation and information retrieval, and the synthesis and general generation of code from text. [47] adds further applications of AI assistance such as code completion, API migration and code repair.
Furthermore, [3] presents "[t]he naturalness hypothesis[:] Software is a form of human communication; software corpora have similar statistical properties to natural language corpora; and these properties can be exploited to build better software engineering tools.". Given the intrinsic formats and formal characteristics of coding and modelling languages employed in SE, the second aspect of the naturalness hypothesis can be considered relatively straightforward. Nevertheless, the initial proposition reinforces the necessity for SE artefacts that are readily comprehensible. This assertion was previously made by [33] in a different form: "[A]ny fool can write code that a computer can understand, good programmers write code that humans can understand.". Consequently, the application of AI in SE is not merely concerned with code generation or code repair; it also encompasses the enhancement of code through techniques such as refactorings, with the objective of optimising readability and maintainability.
Figure 3: AI-assisted Coding
So, while a vast amount of approaches to utilising AI in SE have emerged in the scientific literature, they can be broadly categorised into three principal lines of enquiry: supporting the understanding of SE artefacts, the generation of SE artefacts from another, and the improvement of SE artefacts. These principal approaches may be applied to a variety of SE artefacts, including, but not limited to, requirements, designs, codes, tests, builds and/or miscellaneous of SE processes.
Figure 4: AI4SE - The AI4SE Taxonomy of AI-assisted Software Engineering
The AI4SE taxonomy was inspired by numerous discussions concerning the application of AI in the field of SE. In the absence of a taxonomy that addressed the various options and intricacies of SE and/or AI, a new taxonomy was developed, drawing inspiration from other surveys (see section 6). The dimensions of the first level of AI4SE are categorised as follows: (1) the purpose of using AI for software engineering, (2) the target of using AI within software engineering, (3) the type of AI used, and (4) the level of autonomy/automation achieved by the AI.
The purpose dimension has the three facets of (1) understanding to gain insight, (2) generating to derive (parts of), and (3) improving to advance SE artefacts and/or SE processing. An approach may address multiple purposes. If it is hybrid in this sense, all its purposes are made explicit, e.g. to support understanding only is denoted by ’1’, or to support both understanding and generation is denoted by ’1, 2’.
The AI4SE application targets consist of the main SE activities as described by SWEBOK [73] and DevOps [51]: (1) Software development includes requirements engineering, architecture and design, coding, testing, and maintenance. (2) Software operations includes deployment, execution, and monitoring. Together with the (3) SE process including management and adjustment, the three form the target of application dimension in the taxonomy. Each of the target options can be performed with the support of models, based on models or even driven by models, or without them. For the sake of comprehension of the taxonomy, the model option is optional for each application target: If models are used, i.e. the application target is MDSE, it is marked with "(m)". Thus, the application target "1.m" refers to software development with models, and "1" to one without. Or, in a more fine-grained analysis, "1.d.m" would be used for model-based testing and "1.d" for testing without models. An approach may address several application targets. If it is hybrid in this sense, all of its goals can be made explicit, e.g. support for coding only is denoted by "1.c", support for both coding and testing is denoted by "1.c, 1.d", or support for coding and model-based testing is denoted by "1.c, 1.d.m". For simplicity, "h" can also be used to indicate a hybrid approach of several targets.
The types of AI are divided into (1) symbolic AI, i.e. knowledge representation, rule-based systems, and logical reasoning, (2) statistical AI, i.e. probabilistic modelling and inference, Bayesian networks, Markov models, Gaussian processes, etc., (3) classical machine learning, i.e. supervised, unsupervised, and reinforcement learning, feature engineering and selection, etc., (4) deep learning, i.e. neural networks, transformers, transfer learning, self-supervised learning, attention mechanisms, etc., (5) generative AI (GenAI), i.e. large language models (LLM), generative adversarial networks, diffusion models, variational autoencoders, autoregressive models, contrastive learning, reinforcement learning from human feedback, etc., (6) agentic AI, i.e. multi-agent systems, planning and decision making, hierarchical reinforcement learning, etc., and (7) general purpose AI, i.e. meta-learning and world models. Today, all types (1) to (6) are used in software engineering, but not (7) due to its non-existence. An approach can also use a combination of AI types, i.e. it can be hybrid. This can be made explicit, e.g. by denoting "1, 2" for a hybrid approach that uses both symbolic and classical machine learning. For simplicity, "h" can just be used to denote a hybrid approach that uses more than one type of AI.
We define the levels of autonomy/automation similarly to the levels of autonomous driving [9] as five levels from no support to recommendations to full automation: (1) none, (2) AI-assisted options, where the developer is in full control and receives recommendations to choose from, (3) AI-assisted selection, where the AI pre-selects options, (4) AI-based partial automation, where the AI selects options in simple, standard cases, and (5) AI-based full automation, where the AI operates without the developer. So far, level 3 is the most common, level 4 is on the rise and level 5 is rather unrealistic for complex, industrial-scale software. Higher levels of autonomy typically include the capabilities of lower levels of autonomy; when presenting the level of autonomy of an approach, the highest level of autonomy is given.
Paper Summary Purpose Target Type Level
[55] T-BERT generates trace links between source code and natural language artefacts. It represents a significant advancement in automated software traceability, offering improved accuracy, efficiency, and practicality. 2, 3 1 4 4
[49] NLP approach for improving requirements engineering in the automotive industry: It identifies flaws, such as ambiguous words, incomplete process descriptions, or incorrect quantifiers. 3 1.a 1 3
[63] CBRank method for prioritising software requirements: It uses stakeholder input and requirement attributes to generate an optimal ranking and enables adaptive elicitation by refining the prioritisation based on real-time feedback. 3 1.a, 2.a, 3.1 1, 3 3
[12] ADeX is a tool designed to automate the curation of architectural design decisions in software development. It helps automate the management of architectural decisions, reducing documentation effort while improving decision accuracy. 1, 3 1.b, 2.a 1, 3 4
[14] Explores the early experiences of developers using the AI-augmented coding assistant GitHub Copilot. It may may expand to debugging, code review, and software maintenance, requiring better trust and provenance tracking. 1, 2, 3 1.c 5 4
[78] IntelliCode Compose is a code completion tool that uses a generative transformer model to predict sequences of code tokens, including entire lines of code. It has been trained on a massive dataset of 1.2 billion lines of code across multiple programming languages. 2, 3 1.c 2, 4, 5 4
[6] Getafix is an automated bug-fixing tool that learns from past, human-written fixes to suggest human-like fixes for static analysis warnings. 2, 3 1.c, 1.d 2, 3 3
[38] GraphCodeBERT improves code understanding by incorporating dataflow graphs into the BERT transformer architecture. Data flow graphs represent dependencies between code variables and improve semantic understanding over treating code as a simple sequence of tokens. 1, 3 1.c, 1.d, 1.e 2, 4, 5 4
[40] DeepCodeReviewer is an automatic code analysis system that learns from historical peer reviews to identify and suggest relevant reviews for code snippets. 1, 3 1.c, 1.d, 3.a 2, 3, 4 4
[28] Introduces DeepDebug or automated bug detection and fixing in Java: It generates non-deletion fixes that modify instead of just deleting code 3 1.c, 1.e 4, 5 4
[81] Approach improves code review automation with DL by working with pre-trained models on raw source code and English text. 3 1.c, 1.d 5 3
[79] The book discusses the application of AI beside others for infrastructure monitoring and anomaly detection, predictive maintenance, and build pipeline optimization. 1, 3 1.c, 1.d, 2.a, 3.b 2, 3 4
[16] Presents RepairAgent for automated program repair that allows to autonomously plan and execute actions, such as gathering information, searching for code snippets, and validating fixes, by interacting with a set of specifically designed tools. 3 1.c, 2.a 5, 6 4
[52] Approach links different types of testing information to help automate the generation of equivalence classes for testing. The linking supports tasks such as test reduction, prioritisation and selection. 1, 2, 3 1.d 2, 3 3
[87] Presents a hybrid approach to test case prioritization for regression testing aiming to minimise testing costs. Aims to enhance software testing efficiency by optimising test case execution order. 3 1.d, 1.e 1, 2, 3 3
[75] Introduces Retecs for automatic test case prioritisation and selection in continuous integration. It uses test case meta-data to adapt to changes in the environment and prioritise error-prone test cases. 3 1.d, 2.a 3, 4 4
[21] Explores machine learning-based ensemble methods to predict software quality more accurately than previous approaches 1, 3 1.d, 3.a 3 4
Table 1: AI-assisted software engineering (sorted by application targets)
This novel AI4SE taxonomy [^2] If you want to work on it further, please contact the author for the ontology file which further refines the targets of applications and the AI types. is shown in Figure 4. The fidelity of this new taxonomy AI4SE is verified by reviewing and classifying highly cited publications, but also recent publications to demonstrate the applicability of this taxonomy. The classification of the papers is given in table 1. Additional details on the papers that target mainly the software development are given here:
For requirements, [49] presents an AI-based automation approach to requirements engineering that begins by converting natural language into an Eclipse Modelling Framework (EMF) model. It then applies linguistic rules to identify errors, such as ambiguities or incorrect quantifiers, and provides suggestions for requirements analysts to make final decisions. This approach supports the entire requirements elicitation and change process. Another requirements improvement is given in [63] presenting the prioritisation of requirements by combining the preferences of project stakeholders with approximations of the order of requirements computed by ML techniques.
For architecture & design, [12] describes the automated curation of design decisions to support architectural decision-making. It helps software architects by organizing and recommending design decisions based on previous cases and contextual information of the current project. By leveraging existing design knowledge, the approach analyses historical data and design choices to improve the quality and consistency of architectural designs.
The list of publications on AI-assisted coding, testing, and maintenance is huge. Major developments are for example described in [40] for code understanding. It introduces DeepCodeReviewer that uses deep learning to recommend code reviews for common issues based on historical peer reviews. It assess the relevance of reviews to specific code snippets, suggests appropriate reviews from a repository of common feedback, and improves code reviews by focusing on defect detection. [38] describes GraphCodeBERT, a pre-trained model for programming languages that incorporates data flow semantics rather than just code syntax. GraphCodeBERT demonstrates its performance both in code understanding for code search and clone detection, and in code generation and improvement through code translation and code refinement. [21] presents a study focused on predicting software quality using defect density as a key feature representing quality to achieve higher accuracy in software quality prediction compared to previous studies. The research shows that data pre-processing, feature extraction and the application of ML algorithms significantly improve prediction accuracy. [14] discusses early experiences of developers using GitHub Copilot, which uses a language model trained on source code. Guided by Copilot, developers can write code faster than a human colleague, potentially accelerating development. Three empirical studies with Copilot highlight the different ways developers use Copilot, the challenges they face, the evolving role of code review, and the potential impact of pair programming with AI on software development. [78] discusses IntelliCode Compose, a multilingual code completion tool that predicts entire sequences of code tokens up to full lines of code. The generative transformer model has been trained on 1.2 billion lines of Python, C#, JavaScript and TypeScript code. [6] presents Getafix, a tool for fixing common bugs by learning from previous human-written fixes. It uses hierarchical clustering to group bug fix patterns into a hierarchy from general to specific, and a ranking system based on the context of the code change to suggest the most appropriate fix. Another debugging approach, DeepDebug, is presented in [28], which has been trained by mining GitHub repositories to detect and fix bugs in Java methods.
Since different software testing techniques are complementary, reveal different types of defects and test different aspects of a program, [52] presents an ML-based approach to link test results from different techniques, to cluster test data based on functional similarities, and to generate classifiers according to test objectives, which can be used for test case selection and prioritisation. [81] discusses a test generation approach for writing unit test cases by generating assert statements. The approach uses a transformer model that was first pre-trained on an English text corpus, further semi-supervisedly trained on a large source code corpus, and finally fine-tuned for the task of generating assert statements for unit tests. The assert statements are accurate and increase test coverage. [87] discusses test case prioritisation based on source code changes, software quality metrics, test coverage data, and code coverage-based clustering. It reduces the impact of similar test cases covering the same code and improves fault detection performance.
For deployments, [79] explores the role of ML in DevOps for intelligent release management. It suggests besides other things combining continuous monitoring, predictions of the likelihood of deployment failures, root cause analysis, and pipeline optimisation to reduce deployment failures and improve release management efficiency and software quality.
For the SE process, [75] introduces Retecs, a method for automatically learning test case selection and prioritisation in continuous integration, aimed at minimising the time between code commits and developer feedback on failed tests. Retecs uses reinforcement learning to select and prioritise test cases based on their execution time, previous execution history and failure rates. It effectively learns to prioritise error-prone test cases by following a reward function and analysing past CI cycles. [55] presents the T-BERT framework for generating trace links between source code and natural language artefacts such as requirements or code issues. It demonstrates superior accuracy and efficiency for software traceability, especially in data-limited environments.
Overall, the majority of the presented selection of recent research focuses on improving the software engineering process, either by improving coding, requirements engineering, testing, or the overall processes. A smaller number of approaches aim to improve the understanding of software engineering artefacts. There are also methods that focus on generating code or code-related content. A mixture of AI types are used, including symbolic, statistical, ML, DL and GenAI, as well as hybrid approaches. Many of the AI approaches offer AI-assisted selection, i.e. they provide suggestions, recommendations and rankings that can help developers, but require a developer to make the final decision on what to do. AI-based partial automation is also a common level, where tools automate part of the process, but may require a human to review or further integrate the results. The facets of the AI4SE taxonomy are adequately addressed, although not all options are covered. Nevertheless, it is only a matter of time before research results are presented in primary studies for the open combinations of the facets, as long as they provide advances beyond the state of the art.
5 AI-Assisted Model-Driven Software Engineering
The substantial increase in the utilisation of AI in SE is in parallel accompanied by a notable rise in the application of AI to MDSE. This phenomenon can also be explained by our novel modelling hypothesis, which is an extension of the naturalness hypothesis [3] to modelling artefacts in software engineering:
The modelling hypothesis
Software is (also) a formalized communication between humans and computers; model corpora, software corpora, and natural language corpora have similar statistical properties; the properties of model corpora can be employed to develop more efficacious software engineering tools.
Indeed, the advent of big code and AI has opened up new avenues for integrating AI into model-driven approaches, as posited by the novel modelling hypothesis. As demonstrated by [42], the intrinsic formal graph structures of SE models can also be leveraged to facilitate the preparation of models for gaining AI advantages in MDSE. This is analogous to the considerations in [64] on the necessity to elevate abstraction levels in language models to enhance the efficacy of AI-assisted automated coding.
Paper Summary Purpose Target Type Level
[50] Discusses NLP-driven model generation for more efficient model construction. Presents "purposive" meta-models that guide the interactions between domain experts and GenAI to generate effective prompts for model construction. 2 1.a.m, 1.b.m 1, 5 4
[22] Discusses the NLP-driven domain modelling by evaluating different LLMs and various prompt engineering techniques. 2 1.a.m, 1.b.m 5 4
[80] Details the creation and validation of a high quality dataset of UML diagrams. Provides a valuable research resource in the form of the UML dataset and presents a new, improved approach to UML diagram classification. 3 1.a.m, 1.b.m, 1.c,m, 1.d.m 4 3
[72] Presents a deep learning approach for automatically classifying images of UML diagrams. 1 1.b.m 4 3
[8] Presents an approach to automatically generate model transformations from examples. The approach is demonstrated through experimental evaluation on seven diverse transformation problems. 1, 2 1.b.m 1, 2, 3 4
[57] Explores the automated classification of UML class diagrams into forward-engineered and reverse-engineered diagrams. It provides a dataset with ground truth for the classification of class diagrams. 1, 3 1.b.m 3 3
[1] Presents an experience report detailing the use of LLMs to analyse and transform software variants represented in various formalisms (Java, UML, GraphML, statecharts). 1, 2 1.b.m, 1.c.m 5, 6 3
[29] Presents a software architecture for the AI-augmented automation of DevOps pipelines to address both architectural and technical challenges associated with the development and operation of complex systems. 1 1.b.m, 1.e.m 4 3
[5] Proposes a approach for statistically analysing large collections of models within the field of MDSE. The approach demonstrates its effectiveness in domain analysis and repository management, and shows the potential for applications such as model clone detection. 1 1.b.m, 1.e.m 2, 3 2
[54] Introduces FlowGen, a code generation framework that leverages multiple LLM agents to emulate different software process models (Waterfall, TDD, and Scrum). On standard benchmarks, the Scrum-based FlowGenScrum consistently outperforms other models and baselines. 2 1.c.m, 1.d.m, 1.e.m, 3.1.m, 3.2.m 5, 6 4
Table 2: AI-augmented model-driven software engineering (sorted by application targets)
As shown in Table 2, the growing trend in the adoption of AI for MDSE is reflected by various new research results:
Foundational work on ground truths for developing and benchmarking AI applications in MDSE is discussed in [80, 72] and [57]. [80] presents the creation and validation of a new, high-quality ground truth dataset of UML diagrams. The curated dataset of 2626 unique and accurately labelled images across six UML diagram types goes beyond the Lindholmen dataset of modelling artefacts [66, 67]. The new ground truth dataset was also presented for an improved approach to UML diagram classification. Also [72] presents a new dataset of images, manually labelled into four UML diagram types (class, activity, sequence, use case) and a non-UML category, which is used to develop new DL approaches with higher precision to automate UML diagram classification. Another ground truth dataset is described in [57]: A manually labelled subset of the Lindholmen dataset is used to explore the automated classification of UML class diagrams into forward-engineered (FwCD) and reverse-engineered (RECD) diagrams using supervised machine learning. The methodology in [5] uses techniques from information retrieval, NLP, and ML to enable an efficient comparison, clustering and clone detection of model artefacts.
[50] presents a methodology for AI-augmented MDSE as a continuous cycle of AI-generated model proposals by LLMs that are then reviewed and refined by humans. It seeks to bridge the gap between the natural language used by domain experts and the formal languages used in MDSE.
[22] investigates the feasibility of using LLMs to automate domain modelling in SE. It applies various prompt engineering techniques to a dataset of diverse domain modelling examples and finds that while LLMs show promising domain understanding and achieve relatively high precision in generating classes and attributes, their recall remains low, particularly for relationships. It concludes that full automation is currently impractical.
[8] presents a novel method for automatically generating model transformations from examples. It employs a three-step process: (1) analysing example model pairs to identify groups of similar transformations, (2) using genetic programming to learn basic rules for each group, and (3) refining these rules using simulated annealing to handle complex dependencies and value derivations. The method presented improves the efficiency and accuracy of learning complex model transformations.
[1] explores the use of LLMs to aid in the complex process of re-engineering existing software variants into a software product line (SPL) by identifying features to represent variability and commonalities. It describes the methodology used to prompt the LLMs, the results obtained, and the limitations encountered. There is a potential in assisting domain analysts and developers in transforming software variants into SPLs, but LLMs are limited in managing large variants or complex structures.
[29] presents a novel software architecture designed to address the increasing complexity of modern system development within a DevOps framework. The architecture integrates MDSE and AI techniques to enhance the entire software lifecycle, from requirements engineering to monitoring and maintenance. The architecture is evaluated through multiple industrial case studies.
[54] uses agentic AI where each agent embodies a specific development role such as requirement engineer, developer, or tester and collaborates via chain-of-thought prompting. The set of LLM agents emulate different software process models and are evaluated against four benchmarks (HumanEval, HumanEval-ET, MBPP, and MBPP-ET). The impact of individual development activities on code quality is also investigated, revealing the critical role of testing and the beneficial effects of design and code review.
All papers share the common view that the strengths of AI and MDSE should be combined to further improve the quality of software products and/or processes. While acknowledging the potential of AI-supported MDSE, the papers also highlight the limitations of LLMs, their sensitivity to prompt variations, their limited ability to handle complex inputs, and the risk of inaccuracies. There is a consensus that a fully automated approach is not yet feasible and that human review remains necessary to counteract inaccuracies introduced by AI. UML is seen in these papers as a common modelling language, both for creating models and as an input to AI models. There is also an expectation that AI techniques will support better model quality and improve software reuse by identifying and adapting components from previous projects. This will be supported by the use of model matching and clone detection techniques.
So when AI and MDSE are combined, it becomes clear that there are significant opportunities:
  1. (1)
    The provision of large amounts of top-down models on coding platforms.
  2. (2)
    The generation of bottom-up models from big code.
  3. (3)
    The formation of Big Models as a basis for more advanced empirical MDSE and further improvements of MDSE.
  4. (4)
    The application of AI methods on Big Models and for Big Models.
  5. (5)
    The shift towards the new paradigm of pair modelling in SE, which altogether will turn SE into the next generation of software engineering.
Regarding opportunity (1), [77] presented the first entities of the SEMI Software Engineering Models Index, a catalogue of model repositories, and invited further contributions to SEMI. [43] used another approach of mining GitHub for projects including UML models, which could well be combined with the SEMI initiative. As a result of numerous efforts towards SE model collections, the Lindholmen dataset was developed [66] and reviewed in [67]. For recent work on ground truth datasets see table 2.
In view of opportunity (2), extracting representational models from code is in fact the opposite of using models to generate code more efficiently. Since the design, specification, and maintenance of such models can be complex and time-consuming, various techniques have been developed to extract models from code and/or execution traces. These techniques include the extraction of all types of models including for example the extraction of information models, see e.g. [20, 60], of structural models, see e.g. [48, 39], and of behavioural models, see e.g. [23, 56]. The process of extracting models from code and/or execution traces offers the advantage of retrieving models that are up-to-date with the code/traces, but it also carries the potential disadvantage of mismatch with the model representation and abstraction requirements.
With regard to opportunity (3), there are initial studies on Big Models such as [44] which explores the increasing role of modelling, especially in safety-critical software development. It surveyed a range of projects utilising the UML and identified collaboration as the primary rationale for employing models. This is because models facilitate communication and planning for joint implementation efforts within teams, including those who are not directly involved in modelling or are new to the team.
For opportunity (4), the application of AI methods to Big Models is elaborated in first studies as elaborated in table 2 above.
Last but not least for opportunity (5), the pair modelling paradigm naturally evolves from the pair programming paradigm (see e.g. [13]), which has evolved into one of the killer applications of AI in software engineering, where AI-powered tools/agents act as pair programmers and partners in software development (see e.g. [25]).
Based on Big Models and large SE models models for ML, i.e. large models for ML trained with SE models, pair modelling will emerge as a new model-driven software development technique in which a software engineer and an AI tool/agent collaborate in software development. The engineer or tool/agent in the role of the driver writes or improves software artefacts, including models, while the other, in the role of the observer, reviews each element of the software artefact as it is typed into an artefact. The one in the role of observer is also in the role of navigator, considering systemic and strategic aspects of software development: The navigator identifies potential improvements and possible upcoming problems that need to be addressed later on if they are not to be avoided. This allows the driver to focus on the development aspects without losing sight of the cross-cutting and overarching aspects of software development. Like in pair programming in pair modelling, the observer is used as an assurance of the overall resulting software quality and as a guide to high quality software engineering. The two, the driver and observer/navigator, can switch roles. Indeed, the exact interplay of driver and observer/navigator in pair modelling depends on the capabilities of the AI tool and the level of support according to the AI4SE taxonomy, it can provide to the software engineer being in either role.
And as AI is increasingly applied to Big Models, it remains to be seen whether it becomes necessary to include more specifics of modelling in the AI4SE taxonomy, which currently considers modelling as an optional facet in software development, operations and SE processes. At this early stage of development, it is not clear whether modelling needs to be a separate facet or can remain a cross-cutting concern affecting numerous facets within the AI4SE taxonomy.
6 Related Work
The growing intersection of artificial intelligence (AI) and software engineering (SE) represents a highly active field in research and tool development. The sheer volume of research results has necessitated the creation of surveys [2, 86, 30, 3, 31, 65, 32]that review, sort and structure the state of research in this field. Natural language processing (NLP) techniques, as well as (semi-)formal language processing on large code bases (Big Code) for AI-augmented software engineering, are used to address areas such as code generation, defect detection, code repair, code documentation, software testing, software quality and the challenges related to software security and piracy. Another focus is the application of AI in MDSE [18, 26], examining their use in automating tasks like model generation, transformation, and management, alongside considerations for integration and potential challenges like hallucinations.
It is widely acknowledged that artificial intelligence, and machine learning in particular, is becoming increasingly significant in the field of software engineering. The papers emphasise the potential of AI to automate tasks, enhance code quality, and support developers. It is further noted that AI has the potential to relieve developers from repetitive work, thereby allowing them to focus on higher-level design and problem-solving. Additionally, there is a discussion of the potential of LLMs for code completion, code generation, and other tasks, as well as the challenges and shortcomings of current LLMs. A recurring theme in the literature is the call for more empirical data to support the advancement of this field. Hence, the necessity for a data-driven approach to SE is emphasised by several papers. Also, some papers express concerns about potential biases in AI systems that could have an impact on SE and the resulting software quality.
The papers also exhibit different foci: due to the limited dissemination of MDSE, the majority focusses on the use of AI directly on code, using extensive code bases to train models for code generation and analysis. A smaller portion analyses the application of AI within the context of MDSE, exploring how AI can automate model transformations and facilitate model creation. Furthermore, different taxonomies and classifications are discussed: The papers propose different ways to classify the application of AI in SE, which demonstrates that the field is developing and still lacks a unified framework where AI4SE aims at becoming it. Additionally, there is a divergence of opinion regarding the desired level of automation. Some papers advocate for the utilisation of AI as a tool to assist developers, while others envision the development of fully autonomous systems capable of making decisions without human intervention. Generally, the papers are focused on automating repetitive tasks in software development and reducing errors that may arise from human actions.
More specifically, [32] was one of the first to bridge the gap between AI and SE arguing for their fusion to meet emerging software needs. It explores the potential of integrating AI and SE techniques and tools to create more sophisticated and user-friendly software. [65] discusses how the traditionally separate disciplines of AI and SE are increasingly intertwined. It examines the use of AI techniques to improve SE processes (e.g., project management) and the use of SE methods to develop and manage AI systems (e.g., agent-oriented software engineering). [3] introduces the "naturalness hypothesis" (see section 4) and uses it to justify a data-driven approach to SE. It presents a taxonomy of probabilistic models (code generation, representation, and pattern mining models) and describes their underlying assumptions and applications to SE tasks like code completion, bug detection, and program synthesis. The paper identifies key challenges, such as data sparsity and the need to bridge symbolic programming language representations with ML representations. [86] updates the use of the "naturalness hypothesis" for code generation, completion, translation, refinement, summarisation, defect and clone detection by GitHub Copilot and AlphaCode. It also discusses the challenges and opportunities, including computational cost, quality measurement, software security, piracy concerns, and integration with existing development tools such as Xcode. [2] examines current AI applications across the software development lifecycle, from requirements analysis to maintenance, using tools such as GitHub Copilot and AlphaCode. It analyses the benefits (such as increased productivity and improved code quality) and challenges (such as data quality, ethical concerns and skills gaps) of AI adoption in SE. It also aims to provide a holistic framework for integrating AI responsibly, balancing automation with human oversight, and offers actionable recommendations for practitioners.
Wrt. MDSE, [18] explores the history and future of automation within MDSE. It focuses on the transformative potential of AI for MDSE, addresses challenges such as data scarcity and human-computer interaction, and aims to stimulate further research into the automation capabilities of MDSE through AI and extending its application to different domains and new computing paradigms. [26] examines the burgeoning intersection of LLMs and MDSE, focusing on the automation of tasks such as model repository classification and recommender system development. It explores how LLMs can improve the management, exploration and evolution of entire modelling ecosystems, thereby transforming software engineering practices.
As far as taxonomies for SE and AI are concerned, only two have been made explicit: [31] presents the AI-SEAL taxonomy based on three key facets: point of application (process, product, or runtime), five types of AI technology used (e.g. AI tribes such as symbolist or connectionist), and ten levels of automation (e.g. from "human considers alternatives, makes and implements decision" to "computer makes and implements decision if it feels it should, and informs human only if it feels this is warranted"). In AI4SE, we expand the point/target of the application by detailing the activities involved, we expand the types of AI by adopting a more natural classification to include, for example, agentic AI, and we limit the levels of automation to the more established levels of autonomy when discussing AI applications. In addition, we add the purpose of the AI application, which guides the design, configuration and evaluation of an AI application in SE.
[30] explores the current and future potential of software development bots ("DevBots") as a form of agentic AI. It classifies DevBots according to their purpose, initiation, communication methods, and level of intelligence. It contrasts current DevBots with a vision of an "ideal" DevBot, which would be autonomous, adaptive, and possess both technical and social skills, acting more like artificial team members than mere tools. Like AI4SE, the DevBots taxonomy uses a facet approach to classification, but is restricted to bots, which limits its scope. It also uses a facet for purpose (although this is different from the purpose in AI4SE) and a facet for intelligence (which is limited to only two levels of autonomy/automation). It has two other bot-specific facets for bot initiation and communication.
7 Summary
It is anticipated that the integration of AI will enhance the accessibility of software development to a more heterogeneous user base. By automating tasks, AI can free up developers to focus on higher-level design and problem-solving, leading to a more enjoyable and creative development process. Furthermore, it is regarded as the road towards AI-augmented MDSE bridging the gap between higher-level models and concrete code. However, it is crucial to acknowledge the necessity of addressing the ethical implications of AI in SE, particularly in the context of bias, fairness, transparency, and privacy when developing and deploying AI-augmented software tools. The utilisation of pair programming and/or pair modelling with AI has been identified as a catalyst for enhanced collaborative endeavours between humans and AI. In order to facilitate a more extensive integration of AI into tools, a number of challenges must be surmounted. These include the establishment of interoperability among tools, the creation of standards, and the provision of user-friendly coding and modelling tools. In the future, AI-augmented software lifecycle frameworks will (1) have repetitive and often time-consuming tasks being automated, (2) provide intelligent decision-making by analysing Big Code and/or Big Models for deeper insights, (3) provide enhanced quality of software products by detecting requirements inconsistencies. design flaws, software defects, or other issues, early in the development lifecycle, (4) facilitate management of the increasing complexity of modern systems by providing analysis, monitoring and optimisation tools, and (5) enable continuous improvement by feedback loops and adaptive learning, fostering iterative improvement in software products and software processes.
Following a review of the current state of model-driven software engineering practice, this paper examines big code on open-source software platforms and the application of AI in software engineering.The novel taxonomy of AI-assisted software engineering, AI4SE, is developed and used to categorise highly cited and recent publications. Furthermore, the concept of Big Models is defined and explored with a view to future opportunities for further adoption of model-driven software engineering, taking advantage of AI gains. Finally, selected research on AI approaches for MDSE is reviewed and categorised using the AI4SE taxonomy.
The novel AI4SE taxonomy aims to classify and understand the different ways in which AI is used in SE, recognising the growing overlap between AI and SE. Future work will focus on the development of Big Models and AI applications in MDSE and/or detail the other activities in SE.
Acknowledgement
The ideas presented in this paper have been developed through constructive dialogue in the Feldafinger Kreis, the German Testing Board and the Association for Software Quality and Education. The author acknowledges that while authored by her, the writing process was aided by (AI) tools, specifically Google Scholar and ResearchRabbit for determining related work, and NotebookLM, ChatGPT and DeepL for fine-tuning the wording. The author has no competing interests to declare that are relevant to the content of this article.